OpenStudy (anonymous):

Arithmetic Progression: Find the 18th term of a series that has an nth term given by (2 + 3n). Explain.

OpenStudy (anonymous):

Thanks to jim_thompson5910, I can now do this question. 1st term is when n = 1 2nd term is when n = 2 ... 18th term is when n = 18 So an = 2 + 3n --> a18 = 2+3*18 = 56